Solve the system by addition or substitution.
3x – 4y = 8
6x – 2y = 10

Answer by checkley71(8403)

3X-4Y=8 NOW MULTIPLY THIS EQUATION BY -2 & ADD THESE 2 EQUATIONS.
-6X+8Y=-16
6X-2Y=10
-----------------------------
6Y=-6
Y=-6/6
Y=-1 ANSWER.
3X-4*-1=8
3X+4=8
3X=8-4
3X=4
X=4/3 ANSWER.